OpenStudy (anonymous):

Is anyone good at geometry? I really need help/:

OpenStudy (anonymous):

yea I believe I am

OpenStudy (ladiesman217):

what do you need help with

OpenStudy (anonymous):

I need some help with area and circumference stuff, i'm so confused and lost

OpenStudy (ryan123345):

Go ahaid. :)

OpenStudy (ladiesman217):

ok

OpenStudy (ryan123345):

Circumference = \[d \times \pi \]or \[2r \times \pi\]

OpenStudy (anonymous):

okay i have to find the area of a trapezoid for this one..

OpenStudy (anonymous):

isnt area 1/2b+b x h ?

OpenStudy (anonymous):

that's for a triangle

OpenStudy (ryan123345):

yes its a 45-45-90 triangle so the base1 is 17 base2 is 7 height is 5 and you divide all that by 2. I got to go sorry dude. I told you what to do so hope this helped :)

OpenStudy (anonymous):

Okay! no problem thanks (:

OpenStudy (ryan123345):

:) bye. Be on in 2 hours and ill explain it :)

OpenStudy (anonymous):

sounds good ! (: